OpenStudy (anonymous):

What is the distance between points P(2, 7) and Q(10, 13)?

hartnn (hartnn):

same formula! should i post that again ?

OpenStudy (anonymous):

(2-10)^2=64

hartnn (hartnn):

correct! (7-13)^2 =... ?

OpenStudy (anonymous):

36

hartnn (hartnn):

good 64+36 =... ?

OpenStudy (anonymous):

100

hartnn (hartnn):

and \(s=\sqrt{100}=...?\)

OpenStudy (anonymous):

10

hartnn (hartnn):

or rather \(d=\sqrt{100}=...\) d= distance between the points

hartnn (hartnn):

10 is correct :)